OpenStudy (anonymous):

solve the equation by using the quadratic formula . 3x^2-5x=2

OpenStudy (anonymous):

@chycora @countryboy212

OpenStudy (anonymous):

@boricua5 Do you know the equation?

OpenStudy (anonymous):

the quadratic formula equation?

OpenStudy (anonymous):

@boricua5 Yes.

OpenStudy (anonymous):

the quadratic formula is ax^2+bx+c=0

OpenStudy (anonymous):

\[x =\frac{ {-b \pm \sqrt{{b^{2}}-4ac}} }{ 2a }\] For x=0 and the form:\[{0=a^{2}x+bx+c}\] Right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@boricua5 Well, solve for zero and then use the equation.

OpenStudy (anonymous):

i already solve it

OpenStudy (anonymous):

You should get \[3x^{2}-5x-2=0\]

OpenStudy (anonymous):

yes i did